if i have this equation:
y=18x+0

how do i find y and x?

Answer by elima(1433)

You will choose the x and solve for y.
y=18x+0
We will first make x=1;
y=18(1)+0
y=18+0
y=18
Now let's make x=0
y=18(0)+0
y=0+0
y=0
Now make x=-1;
y=18(-1)+0
y=-18+0
y=-18
You want to choose at least 3 so you can make a line on the graph.
=)